40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11902     Entry time: Sat Dec 26 10:34:43 2015
Author: Steve 
Type: Update 
Category: General 
Subject: today 

PMC locked manually and PRM sus damping restored.

Attachment 1: 1d.png  46 kB  | Hide | Hide all
1d.png
ELOG V3.1.3-